# Visual Localization in 3D Maps: Comparing Point Cloud, Mesh, and NeRF Representations

### v1

@@ Abstract
This paper introduces and assesses a cross-modal global visual localization system that can localize camera images within a color 3D map representation built using both visual and lidar sensing. We present three different state-of-the-art methods for creating the color 3D map — namely point clouds, meshes, and neural radiance fields (NeRF). Our system constructs a database of synthetic RGB and depth image pairs from these representations. This database serves as the basis for global localization. We present an automatic approach that builds this database by synthesizing novel images of the scene and exploiting the 3D structure encoded in the different representations. Next, we present a global localization system that relies on the synthetic image database to accurately estimate the 6 Degrees of Freedom (DoF) camera pose of monocular query images. Our localization approach relies on different learning-based global descriptors and feature detectors which enable robust image retrieval and matching despite the domain gap between (real) query camera images and the synthetic database images. We assess the system’s performance through extensive real-world experiments in both indoor and outdoor settings, in order to evaluate the effectiveness of each map representation, the advantages of learning-based features and descriptors, and the benefits against traditional structure-from-motion localization approaches. Our results show that all three map representations can achieve consistent localization success rates of and higher across various environments. NeRF synthesized images show superior performance, localizing query images at an average success rate of . Furthermore, we demonstrate that our synthesized database enables global localization even when the map creation data and the localization sequence are captured when travelling in opposite directions. Our system, operating in real-time on a mobile laptop equipped with a GPU, achieves a processing rate of .
Index Terms:
Localization, Mapping, Sensor Fusion, RGB-D Perception
1Oxford Robotics Inst., Dept. of Engineering Sci., Univ. of Oxford, UK.
2Dept. of Mechanical Engineering, Univ. of Hong Kong, China
I Introduction
@@
Recognizing a previously visited place and estimating a robot/sensor’s accurate metric pose is fundamental to the problem of robot localization. Localization not only addresses the question of determining where a robot is, but it also enables other tasks such as loop detection in Simultaneous Localization and Mapping (SLAM), multi-robot or multi-session mapping, and augmented reality. Robot localization can be achieved using different sensor modalities, with camera [1, 2, 3] and lidar-based approaches [4, 5, 6] the most commonly studied. While cameras offer a compact form factor and are often low-cost, visual localization and mapping systems may struggle with lighting or seasonal changes. Conversely, lidar systems are physically bigger, have higher power consumption and increased cost — factors which pose challenges for their widespread deployment on mobile robots.
@@
For any localization system, a prior representation of the scene—a prior map—is required. The prior map is usually built using the same sensor modality that will be used for later deployment. For example, autonomous industrial inspection has been demonstrated using legged platforms equipped with lidar [7], where a prior map is built by accumulating laser scans, and localization is performed against the map using methods such as ICP [8]. Similarly, purely visual localization methods often build a map using a structure-from-motion pipeline (SfM) such as COLMAP [9], and then localize hierarchically using visual place recognition followed by pose estimation methods such as perspective-n-points (PnP) [1].
@@
Our approach creates a unified map database for each of the three candidate map representations we study. The visual database is made up of RGB and depth images at known poses and is automatically synthesized from color point clouds, meshes, or NeRF maps. These maps are built using vision and lidar data acquired from either industrial-grade Terrestrial Laser Scanners (TLS) or hand-held mapping devices. Using this unified representation, we achieve scalable global localization resilient to scene changes and lighting variations with state-of-the-art visual localization components. An overview of the system operation is shown in Fig. 1.
@@
A versatile and unified approach for global visual localization of a single camera using a database of synthetic RGB and depth images generated from point cloud, mesh, or NeRF 3D color maps.
@@
An extensive evaluation across both indoor and outdoor environments showcasing the effectiveness and robustness of our proposed method.
A performance comparison between the aforementioned multi-modal localization system and two (purely) visual localization systems.
@@ II-A Visual Localization
For localization, visual place recognition is implemented as an image retrieval problem. Methods such as FAB-MAP [13] and DBoW [14] solve the problem using local features, but more recent methods such as NetVLAD [15], PatchNetVLAD [16], EigenPlaces [17] rely entirely on neural architectures to obtain global descriptors for retrieval. Metric pose estimation is then performed via local feature matching using approaches such as SIFT [18], SuperPoint [19], or R2D2 [20]. This type of pipeline has been implemented in methods such as HLoc [1] and [2], which employ a hierarchical approach to large-scale visual localization.
In addition to the aforementioned methods, alternative approaches have been explored in visual localization. MeshLoc [21] shows it is feasible to construct a dense 3D mesh model from multi-view stereo point clouds and to render synthetic images via an OpenGL pipeline to create a visual database. In another study, [22] opts for a different retrieval approach. Rather than using a global descriptor for a query image, they detect instances of buildings in outdoor environments and retrieve the best match from a database of buildings. They show improvement in long-term and large-scale localization datasets over classic hierarchical frameworks.
II-B Lidar Localization
@@
II-C Combined Visual/Lidar Localization
@@
II-D Cross-Modal Localization
For cross-modal localization, most previous work uses a camera to localize in a lidar map. This is the problem we also focus on in this paper. EdgeMap [33] extracts straight lines from the point cloud map to build a 3D edge map, and applies an edge filter for the camera image. Based on a particle filter, the likelihood of each pose hypothesis is calculated through the filtered camera image and the edge map. Building upon the line extraction idea, Yu et al. [34] extract lines from both the camera image and the lidar map. With a predicted pose from a visual odometry (VO) system, the camera pose is iteratively optimized by minimizing projection error based on 2D-3D line correspondence. They demonstrate that their method can greatly reduce drift by registering live images to the lidar map. However, these methods primarily apply to structured environments where lines and edges are prominent features.
Within the context of autonomous driving, researchers have also been developing localization approaches utilizing inexpensive and readily available camera sensors. Wolcott et al. [35] propose a method to generate a greyscale synthetic view from a mesh map through shading. By minimizing the normalized mutual information between live and synthetic images, this method estimates a pose within the lidar prior map. Similarly, Pascoe et al. [36] synthesize color images from a textured mesh and can further compensate for camera calibration changes.
@@ II-D Cross-Modal Localization
III Method
III-A Problem Definition
@@ III-B System Overview
III-C Pose Selection when Rendering a Visual Database
@@
III-D Generating Synthetic Images
@@
III-D1 Point Clouds — using Terrestrial Lidar Scanning
Point clouds are point samples represented within a 3D coordinate system, sampling an object’s or scene’s external surface. Point clouds are commonly produced by technologies such as lidar scanning and TLS. In our study, we do not utilize point clouds derived from a SLAM system (such as FastLIO [39]), we instead use our TLS-generated point clouds as they are also used to generate precise ground truth maps with millimeter-level measurement accuracy. Further details about the TLS scanning process are provided in Sec. IV-C.
To generate RGB and depth images from the point cloud, we utilize the rendering capabilities of Open3D [40], which leverages an OpenGL framework111OpenGL, Khronos Group, https://www.opengl.org/. Optimal renders are achieved by strategically positioning several light sources around the scene to facilitate specular reflections and shininess. Cameras are placed and maneuvered within the point cloud according to the predefined rendering poses (from the previous subsection). For each pose, 3D points are projected into the camera’s field of view using a pinhole projection model, which allows the pixel size of each point to be adjusted.
@@ III-D1 Point Clouds — using Terrestrial Lidar Scanning
III-D2 Textured Meshes — using ImMesh
@@
In our experiments, we use a state-of-the-art mesh reconstruction framework, ImMesh [43], to reconstruct the triangular mesh of scenes captured using lidar scans. ImMesh initiates this process by estimating the pose of the lidar, followed by registering each scan to a global map. This map is partitioned into fixed-sized volumetric voxels. Mesh reconstruction is carried out online using an incremental, voxel-wise meshing algorithm. Initially, all points within a voxel are projected onto an estimated principal plane to reduce dimensionality. Subsequently, the triangular mesh is reconstructed through a series of voxel-wise operations: mesh pull, commit, and push steps. For more detailed information about the mesh construction process, readers are referred to Section VI of [43]. ImMesh is designed to efficiently reconstruct the mesh of large-scale scenes in real time while maintaining high geometric accuracy.
@@
After reconstructing the textured mesh, we generate synthetic images of the captured scene. This process is accomplished by rendering the mesh into RGB and depth images using OpenGL with the predetermined rendering poses (the approach described in III-C).
III-D3 Neural Radiance Fields — using SiLVR
@@
As with traditional visual 3D reconstruction systems such as Multi-view Stereo, it is difficult for NeRF to reconstruct textureless surfaces or locations with limited multi-view input. However, accurate depth is important for retrieval and matching, as described in Sec. III-E. In addition, training a single NeRF for a large-scale scene is difficult due to relative small model size and limited computational hardware.
@@ III-F Implementation Details
IV Experiment Setup and Hardware
IV-A Dataset
@@
IV-B Hardware - Frontier
@@
IV-C Ground Truth Maps and Trajectories
To generate ground truth point cloud maps, we used a (professional grade) Leica RTC360 Terrestrial Laser Scanner (TLS) to scan the Math and Blenheim sites and an (entry-level) Leica BLK360222https://leica-geosystems.com/products/laser-scanners to scan the ORI site. Views of these maps are shown in the left section of Tab I. TLS scans are captured around apart for indoor ORI and to apart for outdoor Math and Blenheim. Note that ultra-dense distance between TLS scans is impractical due to the considerable time required to scan and process point clouds.
@@ IV-C Ground Truth Maps and Trajectories
These point clouds were also used to render the synthetic images for the point cloud map representation, as explained in Sec III-D1. Before rendering, we voxel-filtered the ORI point cloud to resolution and the Math and Blenheim clouds to .
V Experiment Results
V-A Comparison of Rendered Images
@@
V-B Full System Evaluation
@@
In this section, we quantify the localization performance of the three different rendering pipelines, benchmarking these against two entirely vision-based localization systems, as detailed in Tab. I. The systems we compare are HLoc and COLMAP, both based on Structure-from-Motion (SfM).
Tab. II presents a high-level analysis of the suitability of these systems for key robotics tasks. The SfM approaches are non-metric, meaning their pose estimates are not scaled to real-world dimensions. This limitation necessitates other information to be available to allow effective deployment in robotic applications. These methods are also not real-time, as the SfM bundle adjustment process would take a couple of hours to a day, depending on the image or map size.
@@
V-B1 Our Methods – Cloud, Mesh and NeRF
@@
V-B2 HLoc
HLoc [1] is a 6 DoF visual localization toolkit. It employs a hierarchical localization approach that integrates image retrieval and feature-matching techniques. We first configured HLoc to build an SfM model using all database images and then used it to register query images to the SfM model. For an objective comparison, we used the same pre-trained networks as used in our localization method — NetVLAD for image retrieval, SuperPoint for local feature detection, and SuperGlue for feature matching.
@@ V-B2 HLoc
To obtain query image pose estimates (up to the scale metric), we first refine the pose estimates by eliminating any erroneous matches and then use Evo333https://github.com/MichaelGrupp/evo to align the query image poses with the ground truth poses to recover the scaling factor and the alignment transformation. Finally, we apply this scaling factor to all estimated poses. We can then calculate the localization rate.
@@ V-B3 COLMAP
V-C Localization in Directions Unseen during Mapping

### v2

@@ Abstract
Recent advances in mapping techniques have enabled the creation of highly accurate dense 3D maps during robotic missions, such as point clouds, meshes, or NeRF-based representations. These developments present new opportunities for reusing these maps for localization. However, there remains a lack of a unified approach that can operate seamlessly across different map representations. This paper presents and evaluates a global visual localization system capable of localizing a single camera image across various 3D map representations built using both visual and lidar sensing. Our system generates a database by synthesizing novel views of the scene, creating RGB and depth image pairs. Leveraging the precise 3D geometric map, our method automatically defines rendering poses, reducing the number of database images while preserving retrieval performance. To bridge the domain gap between real query camera images and synthetic database images, our approach utilizes learning-based descriptors and feature detectors. We evaluate the system’s performance through extensive real-world experiments conducted in both indoor and outdoor settings, assessing the effectiveness of each map representation and demonstrating its advantages over traditional structure-from-motion (SfM) localization approaches. The results show that all three map representations can achieve consistent localization success rates of and higher across various environments. NeRF synthesized images show superior performance, localizing query images at an average success rate of . Furthermore, we demonstrate an advantage over SfM-based approaches that our synthesized database enables localization in the reverse travel direction which is unseen during the mapping process. Our system, operating in real-time on a mobile laptop equipped with a GPU, achieves a processing rate of .
keywords:
Localization, Mapping, Sensor Fusion, RGB-D Perception1 Introduction
@@
Recognizing a previously visited place and estimating a robot/sensor’s accurate metric pose is fundamental to the problem of robot localization. Localization not only addresses the question of determining where a robot is, but it also enables other tasks such as loop detection in Simultaneous Localization and Mapping (SLAM), multi-robot or multi-session mapping, and augmented reality. Robot localization can be achieved using different sensor modalities, with camera [1, 2, 3] and lidar-based approaches [4, 5, 6] the most commonly studied. While cameras offer a compact form factor and are often low-cost, visual localization and mapping systems often struggle to build large scale accurate 3D maps, especially in real time. Conversely, lidar systems are physically bigger, have higher power consumption and increased cost — factors which pose challenges for their widespread deployment on mobile robots. Because of this, researchers have proposed combining the advantages of both sensors by constructing a 3D dense map using lidar data (potentially aided by cameras), and then performing localization using only the camera [7, 8].
@@
For any localization system, a prior representation of the scene—a prior map—is required. The prior map is usually built using the same sensor modality that will be used for later deployment. For example, autonomous industrial inspection has been demonstrated using legged platforms equipped with lidar [9], where a prior map is built by accumulating laser scans, and localization is performed against the map using of ICP [10]. However, lidar localisation is sensitive to physical scene changes and relies on a good initial estimate for ICP registration. Conversely, purely visual localization methods typically build a map using a structure-from-motion pipeline (SfM) such as COLMAP [11], and then localize within it using visual place recognition followed pose refinement using methods such as perspective-n-points (PnP) [1]. SfM maps could, however, be inflexible as they are constrained to the local descriptors used to build them. The pose estimated achieved by visual localization within a SfM map are not scaled to real-world dimensions. As a result, additional information is required for effective deployment in robotic applications. Furthermore, these methods are generally not real-time as the SfM bundle adjustment process can take anywhere from a few hours to an entire day, depending on the size of the images or the map.
@@
There has been many recent advances in high-quality 3D color mapping — using either mobile SLAM or fixed TLS data. This has creates new avenues for research into techniques which repurpose these maps for robotic localisation. We propose to create a unified synthetic map representation/database that accommodates the different types of maps produced by these new mapping systems. In particular, these 3D mapping systems enable the rendering of novel synthetic images from arbitrary rendering positions. In our approach, the visual localisation database is made up of rendered RGB and depth images at chosen poses and can be automatically synthesized from color point clouds, meshes, or NeRF maps. These 3D color maps can be built using vision and lidar data acquired from either industrial-grade Terrestrial Laser Scanners (TLS) or hand-held mapping devices. Using this unified representation, we achieve scalable global localization in varied environments, resilient to scene changes and lighting variations with visual learning based components. An overview of the system operation is shown in Fig. 1. In addition, we perform a direct comparison with SfM methods to demonstrate the capability of synthesizing views in the opposite direction of travel, as well as highlighting the advantage of reducing the number of required database images. All experiments are conducted in real-world environments, and the corresponding datasets and ground truth maps are available as part of the Oxford Spires Dataset 111https://dynamic.robots.ox.ac.uk/datasets/oxford-spires/.
@@
A versatile and unified approach for global visual localization of a single camera in dense 3D maps, using a database of synthetic RGB and depth images generated from point cloud, mesh, or NeRF representations.
@@
An extensive evaluation across both indoor and outdoor environments comparing visual localisation performance across different map representations.
A performance comparison between the our localization system with synthetic images and two (purely) visual localization systems with real images.
@@ 2.1 Visual Localization
For localization, visual place recognition is implemented as an image retrieval problem. Classic methods such as FAB-MAP [15] and DBoW [16] solve the problem using local features, but more recent methods such as NetVLAD [17], PatchNetVLAD [18], EigenPlaces [19], MixVPR [20] rely entirely on neural architectures to obtain global descriptors for retrieval. Metric pose estimation is then performed via local feature matching using approaches such as SIFT [21], SuperPoint [22], or R2D2 [23]. This type of pipeline has been implemented in methods such as HLoc [1] and [2], which employ a hierarchical approach to large-scale visual localization.
In addition to the aforementioned methods, alternative approaches have been explored in visual localization in more recent studies. MeshLoc [24] shows it is feasible to construct a dense 3D mesh model from multi-view stereo point clouds and to render synthetic images via an OpenGL pipeline to create a visual database. However, a key limitation of MeshLoc is that the retrieval step relies on a database of real images, again limiting the specific location and direction of the localisation. In another study, [25] opts for a different retrieval approach. Rather than using a global descriptor for a query image, they detect instances of buildings in outdoor environments and retrieve the best match from a database of buildings. They show improvement in long-term and large-scale localization datasets over classic hierarchical frameworks.
Works by Trivigno et al. [26] and Zhang et al. [27] employed an iterative strategy to render images and to refine the pose set based on a query image. While this is an effective approach for solving single query images, it requires an initial pose estimate and lacks the real-time capability needed for robotic applications.
2.2 Lidar Localization
@@
2.3 Combined Visual/Lidar Localization
@@
2.4 Cross-Modal Localization
For cross-modal localization, most previous work uses a camera to localize in a lidar map. These works are the most relevant to our paper. EdgeMap [38] extracts straight lines from the point cloud map to build a 3D edge map, and applies an edge filter for the camera image. Based on a particle filter, the likelihood of each pose hypothesis is calculated through the filtered camera image and the edge map. Building upon the line extraction idea, Yu et al. [39] extract lines from both the camera image and the lidar map. With a predicted pose from a visual odometry (VO) system, the camera pose is iteratively optimized by minimizing projection error based on 2D-3D line correspondence. They demonstrate that their method can greatly reduce drift by registering live images to the lidar map. However, these methods primarily apply to structured environments where lines and edges are prominent features.
Within the context of autonomous driving, researchers have also been developing localization approaches utilizing inexpensive and readily available camera sensors. Wolcott et al. [7] propose a method to generate a greyscale synthetic view from a mesh map through shading. By minimizing the normalized mutual information between live and synthetic images, this method estimates a pose within the lidar prior map. Similarly, Pascoe et al. [8] synthesize color images from a color textured mesh and can further compensate for camera calibration changes. Synthetic images produced by this (ten year old) work was rather limited, and localization success was constrained to the vehicle’s path with fixed viewing angles.
@@ 2.4 Cross-Modal Localization
3 Method
3.1 Problem Definition
@@ 3.2 System Overview
3.3 Pose Selection when Rendering a Visual Database
@@
3.4 Generating Synthetic Images
@@
3.4.1 Render Images from Point Clouds
Point clouds are point samples represented within a 3D coordinate system, sampling an object’s or scene’s external surface. Point clouds are commonly produced by technologies such as lidar scanning and TLS. In our study, we do not utilize point clouds derived from a SLAM system (such as FastLIO [42]), we instead use our TLS-generated point clouds as they are also used to generate precise ground truth maps with millimeter-level measurement accuracy. Further details about the TLS scanning process are provided in Sec. 4.3.
To generate RGB and depth images from the point cloud, we utilize the rendering capabilities of Open3D [43], which leverages an OpenGL framework222OpenGL, Khronos Group, https://www.opengl.org/. Optimal renders are achieved by strategically positioning several light sources around the scene to facilitate specular reflections and shininess. Cameras are placed and maneuvered within the point cloud according to the predefined rendering poses (from the previous subsection). For each pose, 3D points are projected into the camera’s field of view using a pinhole projection model, which allows the pixel size of each point to be adjusted.
@@ 3.4.1 Render Images from Point Clouds
3.4.2 Render Images from Textured Meshes
@@
In our experiments, we extend our previous framework on mesh reconstruction, ImMesh [46], to reconstruct the triangular mesh of scenes captured using lidar scans. ImMesh initiates this process by estimating the pose of the lidar, followed by registering each scan to a global map. This map is partitioned into fixed-sized volumetric voxels. Mesh reconstruction is carried out online using an incremental, voxel-wise meshing algorithm. Initially, all points within a voxel are projected onto an estimated principal plane to reduce dimensionality. Subsequently, the triangular mesh is reconstructed through a series of voxel-wise operations: mesh pull, commit, and push steps. For more detailed information about the mesh construction process, readers are referred to Section VI of [46]. ImMesh is designed to efficiently reconstruct the mesh of large-scale scenes in real time while maintaining high geometric accuracy.
@@
After reconstructing the textured mesh, we generate synthetic images of the captured scene. This process is accomplished by rendering the mesh into RGB and depth images using OpenGL with the predetermined rendering poses (the approach described in 3.3).
3.4.3 Render Images from Neural Radiance Fields
@@
As with traditional visual 3D reconstruction systems such as Multi-view Stereo, it is difficult for NeRF to reconstruct textureless surfaces or locations with limited multi-view input. However, accurate depth is important for retrieval and matching, as described in Sec. 3.5. In addition, training a single NeRF for a large-scale scene is difficult due to relative small model size and limited computational hardware.
@@ 3.6 Implementation Details
4 Experiment Setup and Hardware
4.1 Dataset
@@
Again please note that the datasets and ground truth maps used here are part of the Oxford Spires Dataset mentioned in Sec. 1 .
4.2 Hardware - Frontier
@@
4.3 Ground Truth Maps and Trajectories
To generate ground truth point cloud maps, we used a (professional grade) Leica RTC360 Terrestrial Laser Scanner (TLS) to scan the Math and Blenheim sites and an (entry-level) Leica BLK360333https://leica-geosystems.com/products/laser-scanners to scan the ORI site. Views of these maps are shown in the left section of Tab 1. TLS scans are captured around apart for indoor ORI and to apart for outdoor Math and Blenheim. Note that ultra-dense distance between TLS scans is impractical due to the considerable time required to scan and process point clouds.
@@ 4.3 Ground Truth Maps and Trajectories
These point clouds were also used to render the synthetic images for the point cloud map representation, as explained in Sec 3.4.1. Before rendering, we voxel-filtered the ORI point cloud to resolution and the Math and Blenheim clouds to .
5 Experiment Results
5.1 Comparison of Rendered Images
@@
5.2 Full System Evaluation
@@
In this section, we quantify the localization performance of the three different rendering pipelines using these synthesized images, benchmarking performance against two entirely vision-based localization systems (which use real images), as detailed in Tab. 1. The systems we compare are HLoc and COLMAP, both based on Structure-from-Motion (SfM). We selected these two methods for comparison because they are widely adopted in the community and are designed for large-scale real-world reconstructions. Once the SfM step is completed, these methods allow for real-time localization of query images.
Tab. 2 presents a high-level analysis of the suitability of these systems for key robotics tasks. The SfM approaches are non-metric, meaning their pose estimates are not scaled to real-world dimensions. This limitation necessitates other information to be available to allow effective deployment in robotic applications. These methods are also not real-time, as the SfM bundle adjustment process would take a couple of hours to a day, depending on the image or map size.
@@
5.2.1 Our Methods – Cloud, Mesh and NeRF
@@
5.2.2 HLoc
HLoc [1] is a 6 DoF visual localization toolkit. It employs a hierarchical localization approach that integrates image retrieval and feature-matching techniques. We first configured HLoc to build an SfM model using all real images and then used it to register query images to the SfM model. For an objective comparison, we used the same pre-trained networks as used in our localization method — NetVLAD for image retrieval, SuperPoint for local feature detection, and SuperGlue for feature matching.
@@ 5.2.2 HLoc
To obtain query image pose estimates (up to the scale metric), we first refine the pose estimates by eliminating any erroneous matches and then use Evo444https://github.com/MichaelGrupp/evo to align the query image poses with the ground truth poses to recover the scaling factor and the alignment transformation. Finally, we apply this scaling factor to all estimated poses. We can then calculate the localization rate.
@@ 5.2.3 COLMAP
5.3 Image Descriptor, Feature Detector and Matcher
For a fair comparison, we adopted the same image descriptor and feature detector methods as HLoc. We also experimented with image descriptors such as EigenPlaces [19] and MixVPR [20], but NetVLAD marginally outperformed them when tested on real query images and synthetic database images. A more detailed ablation study comparing learning-based and non-learning-based feature detectors and matchers is provided in Sec. 5.6. It is important to note that our approach is not dependent on the specific learning-based detector and matcher — which can be swapped for improved methods as they emerge. The key point here is that by sampling a database of synthetic images, we can leverage dense 3D maps to provide a unified representation across different map types. This approach offers flexibility in generating poses within the 3D map and even allows for an iterative ”render and compare” strategy if needed, while also reducing the number of database images requires, as compared to SfM-based approaches.
5.4 Localization in Directions Unseen during Mapping